# SuperCalo: Calorimeter shower super-resolution

### v2

@@ Abstract
Calorimeter shower simulation is a major bottleneck in the Large Hadron Collider computational pipeline. There have been recent efforts to employ deep-generative surrogate models to overcome this challenge. However, many of best performing models have training and generation times that do not scale well to high-dimensional calorimeter showers. In this work, we introduce SuperCalo, a flow-based super-resolution model, and demonstrate that high-dimensional fine-grained calorimeter showers can be quickly upsampled from coarse-grained showers. This novel approach presents a way to reduce computational cost, memory requirements and generation time associated with fast calorimeter simulation models. Additionally, we show that the showers upsampled by SuperCalo possess a high degree of variation. This allows a large number of high-dimensional calorimeter showers to be upsampled from much fewer coarse showers with high-fidelity, which results in additional reduction in generation time.
@@ I Introduction
To date, most of the existing surrogate modeling approaches have attempted to directly simulate the full-dimensionality calorimeter showers in a single step, i.e. transforming noise $z$ directly to calorimeter showers $x$ [^1] Two exceptions are [27, 34], which generated showers sequentially, layer by layer, conditioned on previous layers. This is quite different than the approach being proposed here, which will upsample coarse-grained showers all-at-once into higher-resolution showers. This direct, one-step approach presents serious challenges for generative models when $x$ is very high-dimensional. In this work we instead explore an alternative paradigm for the first time, which we dub SuperCalo, that applies super-resolution techniques to high-dimensional calorimeter simulations.
Super-resolution methods based on deep learning have shown exceptional promise in image processing tasks, enhancing image quality and reconstructing high-resolution images from low-resolution inputs (see e.g., [35, 36, 37, 38, 39, 40, 41, 42]). However, many of these existing methods generate the high-resolution images in a deterministic way, whereas our problem requires a probabilistic reconstruction of high-dimensional calorimeter showers given lower resolution calorimeter showers.
@@
One motivated ansatz is to say that each coarse voxel is upsampled to its fine voxels with a universal super-resolution function that may be conditioned on some details such as the coarse voxel location and neighboring coarse voxel energies:
@@
In Section II, we describe the details of Dataset 2 of the CaloChallenge. In Section III, we describe two possible choices of coarse voxelization and the resulting SuperCalo setup based on each choice. We also compare the performance of SuperCalo for the two choices of coarse voxelization. Next, we describe how we generate the coarse voxel energies with a flow-based setup in Section IV. In Section V, we use this setup together with SuperCalo to generate the fine voxel energies given the incident energies of the incoming particles. We evaluate the performance of this full model chain by comparing generated and reference distributions, by using classifier-based metrics (as in [14, 15, 23, 27]), and also by performing timing studies. Finally, we conclude in Section VI.
@@ III.1 Coarse voxelization
For each choice of coarse voxelization, we construct a super-resolution flow to generate the relative [^3] As in, normalized to sum to one. fine voxel energies, $\hat{e}_{{\rm fine},i}$, in a coarse voxel given the coarse voxel energy, $E_{\text{coarse},i}$. We recover the fine voxel energies by multiplying them by the associated coarse voxel energy. We refer to the two super-resolution flows as SuperCalo $A$ and SuperCalo $B$.
@@
  • Deposited energy in neighboring coarse voxels in $\alpha$, $r$ and $z$ directions [^4] There is maximum of 6 neighboring coarse voxels for each coarse voxel. For coarse voxels with fewer than 6 adjacent coarse voxels, the missing neighboring coarse voxel energies are padded with zeros.
@@
Note that the fine layer energies are not obtainable from the coarse shower information for choice A. However, they are obtainable from our “Flow-1” (see Table 1 and Section IV) which generates the 45 fine layer energies as an initial step prior to applying SuperCalo. It costs very little extra to have “Flow-1” generate the fine layer energies instead of only the coarse ones. We decided to include the fine layer energies as conditional inputs as doing so enabled SuperCalo $A$ to more accurately model the distribution of fine voxel energies across the layers. [^5] In principle, it is possible to provide only the coarse layer energies as conditional inputs to SuperCalo, should fine layer energies not be available. The dimensions of the conditional inputs and the outputs of SuperCalo are summarized in Table 1.
@@ III.2 Architecture and training
dim of number of layer sizes number of RQS
@@ IV Generating coarse voxel energies with flows
As it is important for the model to capture the inter-layer correlations between fine voxels, we choose SuperCalo A as the super-resolution flow in the full model chain. [^6] We experimented with combining samples from SuperCalo A and SuperCalo B. In particular, the fine voxels from layers 1-5 were generated using SuperCalo $B$, while the other voxels were generated using SuperCalo $A$. While this improved the layer energy $E_{{\rm layer},i}$ distributions for $i=1,2,...,5$, the classifier scores did not improve significantly. The architectures of Flow-1 and Flow-2 are summarized in Table 2. The details of the conditional inputs and the outputs of Flow-1 and Flow-2 are shown in Table 1, while the preprocessing used during training is detailed in Appendix B.
@@ V Full chain results
We generate 100k electron showers using Flow-1, Flow-2 and SuperCalo. The fidelity of the generated showers is determined by comparing against the reference Geant4 samples. As in Sec. III.4, we present the same four layers when comparing layer-level distributions to give a sense of the overall performance across the full depth of the shower.
@@ VI Conclusion
Another promising extension to this work would be to generalize this approach to even higher-dimensional calorimeter showers (e.g., Dataset 3 of the CaloChallenge [57]). This would require a super-resolution flow that generates $\mathcal{O}(100)$ fine voxel energies given a single coarse voxel energy.

### v3

# Calorimeter shower superresolution

@@ Abstract
Calorimeter shower simulation is a major bottleneck in the Large Hadron Collider computational pipeline. There have been recent efforts to employ deep-generative surrogate models to overcome this challenge. However, many of best performing models have training and generation times that do not scale well to high-dimensional calorimeter showers. In this work, we introduce SuperCalo, a flow-based superresolution model, and demonstrate that high-dimensional fine-grained calorimeter showers can be quickly upsampled from coarse-grained showers. This novel approach presents a way to reduce computational cost, memory requirements and generation time associated with fast calorimeter simulation models. Additionally, we show that the showers upsampled by SuperCalo possess a high degree of variation. This allows a large number of high-dimensional calorimeter showers to be upsampled from much fewer coarse showers with high-fidelity, which results in additional reduction in generation time.
@@ I Introduction
To date, most of the existing surrogate modeling approaches have attempted to directly simulate the full-dimensionality calorimeter showers in a single step, i.e. transforming noise $z$ directly to calorimeter showers $x$ [^1] Two exceptions are [27, 34], which generated showers sequentially, layer by layer, conditioned on previous layers. This is quite different than the approach being proposed here, which will upsample coarse-grained showers all-at-once into higher-resolution showers. This direct, one-step approach presents serious challenges for generative models when $x$ is very high-dimensional. In this work we instead explore an alternative paradigm for the first time, which we dub SuperCalo, that applies superresolution techniques to high-dimensional calorimeter simulations.
Superresolution methods based on deep learning have shown exceptional promise in image processing tasks, enhancing image quality and reconstructing high-resolution images from low-resolution inputs (see e.g., [35, 36, 37, 38, 39, 40, 41, 42, 43]). However, many of these existing methods generate the high-resolution images in a deterministic way, whereas our problem requires a probabilistic reconstruction of high-dimensional calorimeter showers given lower resolution calorimeter showers.
@@
One motivated ansatz is to say that each coarse voxel is upsampled to its fine voxels with a universal superresolution function that may be conditioned on some details such as the coarse voxel location and neighboring coarse voxel energies:
@@
In Section II, we describe the details of Dataset 2 of the CaloChallenge. In Section III, we describe two possible choices of coarse voxelization and the resulting SuperCalo setup based on each choice. We also compare the performance of SuperCalo for the two choices of coarse voxelization. Next, we describe how we generate the coarse voxel energies with a flow-based setup in Section IV. In Section V, we use this setup together with SuperCalo to generate the fine voxel energies given the incident energies of the incoming particles. A schematic of this full model chain is shown in Fig. 1. We evaluate the performance of the full model chain by comparing generated and reference distributions, by using classifier-based metrics (as in [14, 15, 23, 27]), and also by performing timing studies. Finally, we conclude in Section VI.
Figure 1: Schematic of full generation chain where the coarse voxel model (comprising of Flow-1 and Flow-2) and SuperCalo are applied sequentially.
@@ III.1 Coarse voxelization
For each choice of coarse voxelization, we construct a superresolution flow to generate the relative [^3] As in, normalized to sum to one. fine voxel energies, $\hat{e}_{{\rm fine},i}$, in a coarse voxel given the coarse voxel energy, $E_{\text{coarse},i}$. We recover the fine voxel energies by multiplying them by the associated coarse voxel energy. We refer to the two superresolution flows as SuperCalo $A$ and SuperCalo $B$.
Note that choices A and B are examples of two-dimensional coarse graining. We also explored three-dimensional coarse grainings but found them to lead to poorer quality results. In particular, the generated samples tend to have the downsides of both choices A and B.
@@
  • Deposited energy in neighboring coarse voxels in $\alpha$, $r$ and $z$ directions
@@
Note that there is maximum of 6 neighboring coarse voxels for each coarse voxel. For coarse voxels with fewer than 6 adjacent coarse voxels, the missing neighboring coarse voxel energies are padded with zeros. Also, the fine layer energies are not obtainable from the coarse shower information for choice A. However, they are obtainable from our “Flow-1” (see Table 1 and Section IV) which generates the 45 fine layer energies as an initial step prior to applying SuperCalo. It costs very little extra to have “Flow-1” generate the fine layer energies instead of only the coarse ones. We decided to include the fine layer energies as conditional inputs as doing so enabled SuperCalo $A$ to more accurately model the distribution of fine voxel energies across the layers. In principle, it is possible to provide only the coarse layer energies as conditional inputs to SuperCalo, should fine layer energies not be available. The dimensions of the conditional inputs and the outputs of SuperCalo are summarized in Table 1.
@@ III.2 Architecture and training
dimension of number of layer sizes number of RQS
@@ IV Generating coarse voxel energies with flows
As it is important for the model to capture the inter-layer correlations between fine voxels, we choose SuperCalo A as the superresolution flow in the full model chain. [^4] We experimented with combining samples from SuperCalo A and SuperCalo B. In particular, the fine voxels from layers 1-5 were generated using SuperCalo $B$, while the other voxels were generated using SuperCalo $A$. While this improved the layer energy $E_{{\rm layer},i}$ distributions for $i=1,2,...,5$, the classifier scores did not improve significantly. The architectures of Flow-1 and Flow-2 are summarized in Table 2. The details of the conditional inputs and the outputs of Flow-1 and Flow-2 are shown in Table 1, while the preprocessing used during training is detailed in Appendix B.
@@ V Full chain results
As illustrated in Fig. 1, we generate 100k electron showers using Flow-1, Flow-2 and SuperCalo. The fidelity of the generated showers is determined by comparing against the reference Geant4 samples. As in Sec. III.4, we present the same four layers when comparing layer-level distributions to give a sense of the overall performance across the full depth of the shower.
@@ VI Conclusion
At the ML4Jets 2023 conference, SuperCalo was presented as one of the best in terms of sample quality for Dataset 2. However, the specific implementation described in this paper suffers in terms of speed compared to the fastest models in the CaloChallenge due to the fact that we are using a MAF (instead of IAF). Nevertheless, it is important to note that the SuperCalo framework is implementable with many other architectures and hence the speed may vary based on the architecture choice.
Another promising extension to this work would be to generalize this approach to even higher-dimensional calorimeter showers (e.g., Dataset 3 of the CaloChallenge [58]). This would require a superresolution flow that generates $\mathcal{O}(100)$ fine voxel energies given a single coarse voxel energy.